OpenStudy (anonymous):

Find the solution of the exponential equation, correct to four decimal places. 3(1 + 10^8x) = 5

OpenStudy (solomonzelman):

\(\Large\color{black}{ 3(1 + 10^{8x}) = 5 }\) \(\Large\color{black}{ (1 + 10^{8x}) = \frac{5}{3} }\) \(\Large\color{black}{ 10^{8x} = 1+\frac{5}{3} }\) \(\Large\color{black}{ 10^{8x} = 1+1\frac{2}{3} }\) \(\Large\color{black}{ 10^{8x} = 2\frac{2}{3} }\) so far so good ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Use the above relation to make your equation with log. \(\Large\color{black}{ A^B=C~~~~~~->~~~~~~~~A^C=B }\) You can do it...

OpenStudy (anonymous):

0.0532?

OpenStudy (anonymous):

jk thats not right